# Radar Anti-jamming Strategy Learning via Domain-knowledge Enhanced Online Convex Optimization

### v2

@@ 1 Introduction
To address the limitations of traditional jamming countermeasures, frequency agility has been adopted, enabling radar systems to implement adaptive frequency hopping strategies in the face of main lobe jamming [5, 6, 7]. Recently, the application of learning techniques has redefined anti-jamming efforts as sequential decision-making problems. In particular, reinforcement learning has been incorporated into the development of anti-jamming strategies, with notable research exploring the deployment of Deep Q-Networks (DQN) [8, 9]. Game-theoretic methods, including two-person zero-sum games and extensive-form games with imperfect information, have been leveraged to enhance strategic anti-jamming performance [10, 11, 12]. Additionally, the multi-armed bandit framework has emerged as a novel approach for designing anti-jamming strategies [13, 14]. The field continues to advance with studies on subpulse-level frequency agile radar systems [15] and the development of adaptive power allocation models for jammers [11]. However, these active online learning techniques in anti-jamming strategy design either fall short in sample efficiency or struggle to model the dynamics of the jammer.
@@ 2.1 Signal Model
Fig. 1: A detailed illustration on the anti-jamming scenario
@@ 3.1 Online Optimization Formulation
Based on the established signal models and post-processing procedure in Section 2, each pulse can be regarded as one round of interaction between the radar and jammer. Protocol 1 presents the protocol for this type of iterative interaction. The radar system dynamically adjusts its transmitted signal $s_{n}(t)$ to improve the SINR, by utilizing information extracted from previous pulses. To systematically select parameters of $s_{n}(t)$ at each iteration for this online interaction process, we propose an online optimization framework from the radar’s perspective. The goal is to enhance the radar’s adaptive capabilities in response to jammer, ensuring that with each pulse, the radar’s response is optimally adjusted to the evolving jamming strategies. Basic elements involved in this online interaction process are given below.
@@ 3.2 Online Mirror Descent
Setting $\nabla f_{n}=\boldsymbol{\hat{\ell}}_{n}$, the OMD algorithm in (9) is equivalent to a classical bandit algorithm known as Exp3 [25], characterized by a sub-linear static regret bound of $\mathcal{O}(\sqrt{N|\mathcal{A}_{\mathrm{R}}|\log{|\mathcal{A}_{\mathrm{R}}|}})$. However, the IWE neglects the strategic nature of the interaction inherent in the online process, where the radar is essentially engaged in a game against a jammer. Critical insights into the jammer’s behavior remain unutilized. By incorporating knowledge about the jammer’s action space and strategy, we can develop more sophisticated and effective online algorithms.
@@ 4.1 Action Modeling
$\mathbb{E}[\boldsymbol{\hat{\ell}}_{n}]=\textbf{U}\times\mathbb{E}[\boldsymbol{b}_{n}]=\textbf{U}\times\boldsymbol{y}_{n}=\boldsymbol{\ell}_{n}.$
$\boldsymbol{b}_{n}$ is the pure strategy of jammer related with its action $b_{n}\sim\boldsymbol{y}_{n}$.
By incorporating the gradient estimator from (12) into the OMD (9), we introduce an enhanced algorithm referred to as OMD with Action Modeling (OMD-AM), which is elaborated in Algorithm 1. OMD-AM achieves a sub-linear static regret bound $\mathcal{O}(\sqrt{N\log{|\mathcal{A}|}})$, with the proof omitted here. This improved regret bound effectively removes the square-root dependence on the size of the action set $|\mathcal{A}|$ that characterizes the classical Exp3 algorithm. The key to this advancement lies in leveraging the inherent game structure of the cost matrix.
Algorithm 1 Online Mirror Descent with Action Modeling(OMD-AM)
@@ 4.2 Opponent Modeling
where $\boldsymbol{\hat{y}}_{n}=\pi_{n}(h_{n})$ represents the estimate of the decision rule $\pi$ at pulse $n$. A natural choice for $\pi_{n}$ is the maximum likelihood estimator (MLE), which calculates the frequency of each action $b\in\mathcal{A}_{\mathrm{J}}$, conditioned on a length-$k$ history $h\in\mathcal{H}$, across the entire interaction history $(a_{1},b_{1},\ldots,a_{n-1},b_{n-1})$. Combining gradient estimator (13) with OMD (9), we attain OMD with Opponent Modeling (OMD-OM), elaborated in Algorithm 2.
Notably, for OMD-OM, the static regret defined in (7) may tend to negative as it only compares with the best strategy $\boldsymbol{x}^{*}$ in hindsight. This comparator is weak since OMD-OM exploits the jammer strategy information. To address this limitation, we introduce a broader performance metric termed universal regret, which aligns more closely with the objective function described in (6).
@@
where $(\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{N})$ is a comparator sequence. Furthermore, if $\boldsymbol{z}_{1}=\ldots=\boldsymbol{z}_{N}=\boldsymbol{x}=\operatorname*{argmin}_{\boldsymbol{x}}\sum_{n=1}^{N}f_{n}(\boldsymbol{x})$, it reverts to static regret. Typically, achieving sub-linear universal regret is unattainable due to the unknown dynamics of the environment. Nevertheless, OMD-OM, which exploits the information of jamming strategy generation, utilizes a history-dependent predictor to capture the dynamic nature of the jamming strategy. By doing so, it can attain a sub-linear universal regret bound, expressed as $\mathcal{O}(\sqrt{N|\mathcal{H}|\log{|\mathcal{A}_{\mathrm{R}}|}})$ with the proof omitted here. This bound represents a significant improvement over the sub-linear static regret bounds obtained with Exp3 and OMD-AM, offering a mechanism to adapt to the dynamics of jamming environments effectively.
Algorithm 2 Online Mirror Descent with Opponent Modeling (OMD-OM)
@@ 5 Experiments
This section evaluates the efficiency of the developed OMD-AM and OMD-OM algorithms through simulation experiments. Detailed settings for the radar and jammer systems are provided in Table 1. The static regret (7) and universal regret (14) are employed as the performance metrics. Two different jamming strategies were simulated: a stationary strategy where $\boldsymbol{y}_{n}=\boldsymbol{y}$ remains constant for all iterations within the simplex $\Delta_{|\mathcal{A}_{J}|}$, and a non-stationary strategy that varies in response to recent length-$3$ history, denoted as $\boldsymbol{y}_{n}=\pi(h_{n}),h_{n}\in\mathcal{H}$. Specifically, $\boldsymbol{y}_{n}(b_{1})=0.7$ and $\boldsymbol{y}_{n}(b_{2})=0.3$ where $b_{1},b_{2}$ are related with the two most common frequencies in $h_{n}$. The Exp3 algorithm serves as the baseline for comparison. For each algorithm, 500 independent trials are conducted, and the shaded areas in Fig.2 represent the confidence intervals of these trials.
@@
Fig. 1(a) compares the average static regret under the stationary jamming strategy. The results indicate that all the evaluated algorithms achieve sub-linear regret. Notably, both OMD-OM and OMD-AM algorithms outperform the baseline Exp3 algorithm and OMD-OM significantly reduces the number of samples.
@@
In Fig. 1(b), both the static and universal regrets for different algorithms under a non-stationary jamming strategy are compared. With regard to static regret, the OMD-AM algorithm exhibits a more rapid decline compared to the baseline Exp3, indicating a faster convergence rate. Notably, the OMD-OM algorithm demonstrates a substantial reduction in static regret, highlighting its smaller cumulative cost compared with the optimal static strategy. For universal regret, only OMD-OM achieves a sub-linear regret, showcasing its impressive capability to adapt to non-stationary environments.
Finally, Tab. 2 compares the instantaneous cost incurred by different algorithms under both stationary and non-stationary jamming strategies, evaluated after $1000$ rounds. The costs are calculated from E.q. (5). This table clearly underscores the superior performance of the OMD-AM and OMD-OM algorithms under different jamming strategies.
(a) Stationary jamming strategy.
(b) Non-stationary jamming strategy.
Fig. 2: Comparison analysis of three algorithms under two different jamming strategies.
Jamming strategy Exp3 OMD-AM OMD-OM
Stationary 0.25 0 0
Non-stationary 0.42 0.34 0.09
Table 2: Instantaneous costs for various algorithms after $1000$ rounds.

### v3

@@ 1 Introduction
To address the limitations of traditional jamming countermeasures, frequency agility has been adopted, enabling radar systems to implement adaptive frequency hopping strategies in the face of main lobe jamming [5, 6, 7]. Recently, the application of learning techniques has redefined anti-jamming efforts as sequential decision-making problems. In particular, reinforcement learning has been incorporated into the development of anti-jamming strategies, with notable research exploring the deployment of Deep Q-Networks (DQN) [8, 9]. Game-theoretic methods, including two-person zero-sum games and extensive-form games with imperfect information, have been leveraged to enhance strategic anti-jamming performance [10, 11, 12]. Additionally, the multi-armed bandit framework has emerged as a novel approach for designing anti-jamming strategies [13]. The field continues to advance with studies on subpulse-level frequency agile radar systems [14] and the development of adaptive power allocation models for jammers [11]. However, these active online learning techniques in anti-jamming strategy design either fall short in sample efficiency or struggle to model the dynamics of the jammer.
@@ 2.1 Signal Model
Fig. 1: Illustration on the anti-jamming scenario.
@@ 3.1 Online Optimization Formulation
Based on the established signal models and post-processing procedure in Section 2, each pulse can be regarded as one round of interaction between the radar and jammer. Protocol 1 presents the protocol for this type of iterative interaction. The radar system dynamically adjusts its transmitted signal $s_{n}(t)$ to improve the SINR, by utilizing information extracted from previous pulses. To systematically select parameters of $s_{n}(t)$ at each iteration for this online interaction process, we propose an online optimization framework from the radar’s perspective. The goal is to enhance the radar’s adaptive capabilities in response to jammer, ensuring that with each pulse, the radar’s response is optimally adjusted to the evolving jamming strategies. Basic elements involved in this interaction process are given below.
@@ 3.2 Online Mirror Descent
Setting $\nabla f_{n}=\boldsymbol{\hat{\ell}}_{n}$, the OMD-IWE algorithm in (9) is equivalent to a classical bandit algorithm known as Exp3 [24], characterized by a sub-linear static regret bound of $\mathcal{O}(\sqrt{N|\mathcal{A}_{\mathrm{R}}|\log{|\mathcal{A}_{\mathrm{R}}|}})$. However, the IWE neglects the strategic nature of the interaction inherent in the online process, where the radar is essentially engaged in a game against a jammer. Critical insights into the jammer’s behavior remain unutilized. By incorporating knowledge about the jammer’s action space and strategy, we can develop more sophisticated and effective online algorithms.
@@ 4.1 Action Modeling
$\mathbb{E}[\boldsymbol{\hat{\ell}}_{n}]=\textbf{U}\times\mathbb{E}[\boldsymbol{b}_{n}]=\textbf{U}\times\boldsymbol{y}_{n}=\boldsymbol{\ell}_{n}$
and $\boldsymbol{b}_{n}$ is the pure strategy of its action $b_{n}\sim\boldsymbol{y}_{n}$.
By incorporating the gradient estimator from (12) into the OMD (9), we introduce an enhanced algorithm referred to as OMD with Action Modeling Estimator (OMD-AME), which is elaborated in Algorithm 1. OMD-AME achieves a sub-linear static regret bound $\mathcal{O}(\sqrt{N\log{|\mathcal{A}|}})$, with the proof omitted here. This improved regret bound effectively removes the square-root dependence on the size of the action set $|\mathcal{A}|$ that characterizes the classical OMD-IWE (Exp3) algorithm. The key to this advancement lies in leveraging the inherent game structure of the cost matrix.
Algorithm 1 Online Mirror Descent with Action Modeling Estimator (OMD-AME)
@@ 4.2 Opponent Modeling
where $\boldsymbol{\hat{y}}_{n}=\pi_{n}(h_{n})$ represents the estimate of the decision rule $\pi$ at pulse $n$. A natural choice for $\pi_{n}$ is the maximum likelihood estimator (MLE), which calculates the frequency of each action $b\in\mathcal{A}_{\mathrm{J}}$, conditioned on a length-$k$ history $h\in\mathcal{H}$, across the entire interaction history $(a_{1},b_{1},\ldots,a_{n-1},b_{n-1})$. Combining gradient estimator (13) with OMD (9), we attain OMD with Opponent Modeling Estimator (OMD-OM), elaborated in Algorithm 2.
Notably, for OMD-OME, the static regret defined in (7) may tend to negative as it only compares with the best strategy $\boldsymbol{x}^{*}$ in hindsight. This comparator is weak since OMD-OME exploits the jammer strategy information. To address this limitation, we introduce a broader performance metric termed universal regret, which aligns more closely with the objective function described in (6).
@@
where $(\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{N})$ is a comparator sequence. Furthermore, if $\boldsymbol{z}_{1}=\ldots=\boldsymbol{z}_{N}=\boldsymbol{x}=\operatorname*{argmin}\limits_{\boldsymbol{x}}\sum\limits_{n=1}^{N}f_{n}(\boldsymbol{x})$, it reverts to static regret. Typically, achieving sub-linear universal regret is unattainable due to the unknown dynamics of the environment. Nevertheless, OMD-OME, which exploits the information of jamming strategy generation, utilizes a history-dependent predictor to capture the dynamic nature of the jamming strategy. By doing so, it can attain a sub-linear universal regret bound, expressed as $\mathcal{O}(\sqrt{N|\mathcal{H}|\log{|\mathcal{A}_{\mathrm{R}}|}})$ with the proof omitted here. This bound represents a significant improvement over the sub-linear static regret bounds obtained with traditional OMD-IWE (Exp3) and improved OMD-AME algorithm, offering a mechanism to adapt to the dynamics of jamming environments effectively.
Algorithm 2 Online Mirror Descent with Opponent Modeling Estimator (OMD-OME)
@@ 5 Experiments
This section evaluates the efficiency of the developed OMD-AME and OMD-OME algorithms through simulation experiments. Detailed settings for the radar and jammer systems are provided in Table 1. The static regret (7) and universal regret (14) are employed as the performance metrics. Two different jamming strategies were simulated: a stationary strategy where $\boldsymbol{y}_{n}=\boldsymbol{y}$ remains constant for all iterations within the simplex $\Delta_{|\mathcal{A}_{J}|}$, and a non-stationary strategy that varies in response to recent length-$3$ history, denoted as $\boldsymbol{y}_{n}=\pi(h_{n}),h_{n}\in\mathcal{H}$. Specifically, $\boldsymbol{y}_{n}(b_{1})=0.7$ and $\boldsymbol{y}_{n}(b_{2})=0.3$ where $b_{1},b_{2}$ are related with the two most common frequencies in $h_{n}$. The OMD-IWE (Exp3) algorithm serves as the baseline for comparisons on static regret and universal regret. For each algorithm, 500 independent trials are conducted, and the shaded areas in Fig.2 represent the confidence intervals of these trials.
@@
Fig. 2(a) compares the average static/universal regret under a stationary jamming strategy. In the stationary environment, the two regrets are identical. The results demonstrate that all evaluated algorithms achieve sub-linear regret. Notably, OMD-OME and OMD-AME outperform the baseline OMD-IWE (Exp3) algorithm, significantly reducing the number of required samples.
@@
In Fig. 2(b) and 2(c), static and universal regrets under a non-stationary jamming strategy are displayed. With regard to static regret in Fig. 2(b), OMD-AME and OMD-OME exhibit more rapid declines compared to the baseline OMD-IWE (Exp3), indicating a faster convergence rate. Notably, OMD-OME demonstrates a substantial reduction in static regret, highlighting its smaller cumulative cost compared with the optimal static strategy. For universal regret, only OMD-OME achieves a sub-linear regret, showcasing its impressive capability to adapt to non-stationary environments.
Finally, Fig. 2(d) compares the normalized SINR values received by the radar system using different online algorithms after $1000$ interactions. This comparison directly demonstrates the effectiveness of each method within a limited sample size. The proposed OMD-AME and OMD-OME achieve superior SINR in both stationary and non-stationary, showcasing their effectiveness and sample efficiency.
Limitations Our study does have certain limitations that should be addressed in future work. A key challenge arises from the practical application scenarios where the set of available frequencies, $\mathcal{F}$ in (1), is typically large. The size of the action set $\mathcal{A}_{\mathrm{R}}$ thus increases exponentially, significantly impairing sample efficiency across all methodologies. Specifically, in our proposed method, although the opponent modeling approach has substantially enhanced sample efficiency, the large action set still demands a more efficient estimation approach for the ’fixed rule’ $\pi$, beyond the traditional MLE method.
(a) Stationary Environment
(b) Non-stationary Environment
(c) Non-stationary Environment
(d) SINR Comparison
Fig. 2: Comparative analysis of baseline OMD-IWE with proposed OMD-AME and OMD-OME methods, focusing on regret and SINR performance. (a) Regret comparison in a stationary env., where static and universal regrets coincide. (b) Static regret comparison in a non-stationary env. (c) Universal regret comparison in a non-stationary env. (d) SINR values after $1000$ interactions across both env.